Division P of Launch Corporation has the capacity for making 81,500 wheel sets per year and regularly sells 66,500 each year on the outside market. The regular sales price is $165 per wheel set, and the variable production cost per unit is $117. Division Q of Launch Corporation currently buys 36,500 wheel sets (of the kind made by Division P) yearly from an outside supplier at a price of $155 per wheel set. If Division Q were to buy the 36,500 wheel sets it needs annually from Division P at $139 per wheel set, the change in annual net operating income for the company as a whole, compared to what it is currently, would be:

Multiple Choice • $665,000 • $667,000 • $1,387,000 • $83,000 • $1,412,725
Answer #1

Div P-

Current scenario = Contribution = 66,500 * $48 = $3,192,000

After change =

Regular Selling Price = $165

Variable costs = $117

Contribution = $48

No of units sold outside = 81500-36500 = 45,000 wheel sets

Outside contribution = 45,000*$48 = $2,160,000

Additional contribution due to Q = $139 - $117 = $22 * 36500 = $803,000

Total contribution = $2,963,000

Loss of contribution after change = $229,000

Div Q -

Purchase price = $155/unit

Price to buy from P = $139/unit

Saving per set = $16

Total savings = 36,500*16 = $584,000

=> Change in net income = 584000-229000 = $355,000

(This option has not been given but this is the answer I am getting)